An alloy needs to contain between 46% copper and 50% copper. Find the least and greatest amount of 60% copper alloy that should be mixed with a 40% copper alloy in order to end up with thirty pounds of an alloy containing an allowable percentage of copper
Answer by nerdybill(7384)

An alloy needs to contain between 46% copper and 50% copper. Find the least and greatest amount of 60% copper alloy that should be mixed with a 40% copper alloy in order to end up with thirty pounds of an alloy containing an allowable percentage of copper
.
Let x = amount of 60% copper alloy
then
30-x = amount of 40% copper alloy
.
Considering least amount:
.60x + .40(30-x) = .46(30)
.60x + 12 - .40x = 13.8
.20x + 12 = 13.8
.20x = 1.8
x = 9 pounds
.
Considering greatest amount:
.60x + .40(30-x) = .50(30)
.60x + 12 - .40x = 15
.20x + 12 = 15
.20x = 3
x = 15 pounds